OpenStudy (anonymous):

write 8 5/7 as a improper fraction

OpenStudy (wolfe8):

Welcome to OpenStudy. Multiply 8 with 7, add it to 5 and that will be your numerator.

OpenStudy (anonymous):

8 5/7 8*7+5= 61/7 is the correct

OpenStudy (solomonzelman):

\[\huge\color{blue}{ a\frac{b}{c} =\frac{(a \times c) + b}{c} } \]

OpenStudy (wolfe8):

Yup, looks right.

OpenStudy (wolfe8):

Welcome again and hope we helped. Do take time to read our CoC: http://openstudy.com/code-of-conduct and enjoy your stay. If you are satisfied with any question, please reward a medal to whoever you think deserves it and close the question. We also have a fan and testimony system. See you around :)